how many joules of heat are released when 110g of water cools from 91 degrees Celsius to 28 degrees celsius


1
Expert's answer
2022-01-06T19:13:02-0500

The specific heat of water "c=4.184\frac{J}{g\cdot\degree{C}}"


"Q=cm\Delta{T}=4.184\frac{J}{g\cdot\degree{C}}\times110\ g\times(91\degree{C}-28\degree{C})=2.9\times10^4\ J"


Answer: "2.9\times10^4\ J"